OpenStudy (anonymous):

sin^4 θ − cos^4 θ = sin^2 θ − cos^2 θ

OpenStudy (anonymous):

I know it's the power lowering thing, but I'm not very good at that

OpenStudy (anonymous):

\[\frac{ 1-\cos2(4x)-1+\cos2(4x) }{ 4 }\]

OpenStudy (anonymous):

is that right so far?

OpenStudy (mertsj):

No. This is a factoring problem. What are the factors of \[x^2-y^2\]

OpenStudy (mertsj):

brandon...are you there?

OpenStudy (anonymous):

yea, I'm trying to figure it out

OpenStudy (anonymous):

2(x-y)?

OpenStudy (mertsj):

\[a^2-b^2=(a-b)(a+b)\]

OpenStudy (mertsj):

Use that to factor x^2-y^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh ok. x^2-y^2=(x-y)(x=Y)

OpenStudy (anonymous):

*(x+y)

OpenStudy (mertsj):

yes. (x+y) Now by extension: \[a^4-b^4=(a^2-b^2)(a^2+b^2)\]

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

So use that to factor \[x^4-y^4\]

OpenStudy (anonymous):

\[x ^{4}+y ^{4}=(x ^{2}-y ^{2})(x ^{2}+y ^{2})\]

OpenStudy (mertsj):

Yes. And do you see that this is the same form: \[\sin ^4\theta-\cos ^4\theta\]

OpenStudy (anonymous):

so it's (sin^2-cox^2)(sin^2+cos^2)

OpenStudy (mertsj):

Yes. And do you know that sin^2 +cos^2=1?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So try your problem again now.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sin ^{4}\theta - \cos^{4}\theta = (\sin^{2}\theta - \cos^{2})(\sin^{2}\theta +\cos^{2}\theta)\]

OpenStudy (mertsj):

So far so good

OpenStudy (anonymous):

ok, what am I looking for next?

OpenStudy (anonymous):

sin^2+cos^2=1, I know, but how to I pull that out of this?

OpenStudy (mertsj):

\[(\sin ^2\theta-\cos ^2\theta )(\sin ^2\theta+\cos ^2\theta)=(\sin ^2\theta-\cos ^2\theta)(1)=\sin ^2\theta-\cos ^2\theta\]

OpenStudy (mertsj):

This is the way the entire problem should look. (I'm going to use x instead of theta): \[\sin ^4x-\cos ^4x=\sin ^2x-\cos ^2x\] \[(\sin ^2x-\cos ^2x)(\sin ^2x+\cos^2x)=\] \[(\sin ^2x-\cos ^2x)(1)=\] \[\sin ^2x-\cos ^2x=\sin ^2x-\cos ^x\]

OpenStudy (anonymous):

ok, cool, I'm going to try to work through that a couple times

OpenStudy (mertsj):

Excellent

OpenStudy (anonymous):

cool, I see how it works. thanks so much!